OpenStudy (anonymous):

HELP ME PLEASE . 1.use a half-angle formula to find the exact value of sin 67.5°. 4. find the exact solution in the interval [0,2pi] for tanx/2=sinx/2 5. rewrite the following expression in terms of the first power if cosine tan^2xsin^2x

OpenStudy (anonymous):

Since we know that: \(\cos(2x)=1-2\sin^2(x)\) We can also say that: \[\cos(\phi)=1-2\sin^2\left(\frac{\phi}{2}\right)\] And we can isolate for sin: \[\sin\frac{\phi}{2}=\pm\sqrt{\frac{1-cos(\phi)}{2}}\] So then we are given the angle 67.5 degrees. Notice how \(67.5=\frac{135}{2}\) So therefore: \[\sin\frac{135}{2}=\pm\sqrt{\frac{1-\cos135}{2}}\]

OpenStudy (anonymous):

Im not done yet haha

OpenStudy (anonymous):

You can find the exact value of \(cos135\) by using a 45-45-90 triangle. It's real easy if you haven't already learned it:  So then If we know that the \(\cos135^\circ=\frac{1}{\sqrt{2}}\) Then we can simplify the answer even more. Just give me a sec

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

K so then: \[\sin\frac{135}{2}=\pm\sqrt{\frac{1-\frac{1}{\sqrt2}}{2}}=\pm\sqrt{\frac{\frac{\sqrt{2}-1}{\sqrt{2}}}{2}}=\pm\sqrt{\frac{\sqrt{2}+1}{2\sqrt{2}}}\]